For the following matrices check the statement: “If a matrix X has full rank then the matrix XX is invertable” and find thea - c
Answer #1

RUR-R2 O Rgt Rg-R, 0 ser Rank (X) =2 full lank xx = 1 2 3 99 30 3 3 11 3 O 4 1 TXIXI 330-1215209 و رعاع -11 30 -11 209 209 -14 Ixlxls 4-I-3 Ixx1= 3 +0 invertible (xxlt يعاملهم Ixixl 1 3 1012 -1 3 नन -1 xx= a. -1/3 413 x = 2 3 3 4 1 R₂ & Ro-R Rgy Įlaak Ronis ex 122. للسه xl: - 0 0 : رابع HT o 0 0 O 1xix1 = 1 to txxjta Sov matrix since xx ů invectible identity mateix 6 i